### v1

# Think in Safety: Unveiling and Mitigating Safety Alignment Collapse in Multimodal Large Reasoning Model

Abstract
The rapid development of multimodal large reasoning models (MLRMs) has demonstrated broad application potential, yet their safety and reliability remain critical concerns that require systematic exploration. To address this gap, we conduct a comprehensive and systematic safety evaluation of 11 MLRMs across 5 benchmarks and unveil prevalent safety degradation phenomena in most advanced models. Moreover, our analysis reveals distinct safety patterns across different benchmarks: significant safety degradation is observed across jailbreak robustness benchmarks, whereas safety-awareness benchmarks demonstrate less pronounced degradation. In particular, a long thought process in some scenarios even enhances safety performance. Therefore, it is a potential approach to addressing safety issues in MLRMs by leveraging the intrinsic reasoning capabilities of the model to detect unsafe intent. To operationalize this insight, we construct a multimodal tuning dataset that incorporates a safety-oriented thought process. Experimental results from fine-tuning existing MLRMs with this dataset effectively enhances the safety on both jailbreak robustness and safety-awareness benchmarks. This study provides a new perspective for developing safe MLRMs. Our dataset is available at https://github.com/xinyuelou/Think-in-Safety.
Warning: this paper contains example data that may be offensive or harmful.
Think in Safety: Unveiling and Mitigating Safety Alignment Collapse in Multimodal Large Reasoning Model
Xinyue Lou1111Equal contribution. Contact e-mail: louxinyue@bjtu.edu .cn, youlee@bjtu.edu.cn, You Li1111Equal contribution. Contact e-mail: louxinyue@bjtu.edu .cn, youlee@bjtu.edu.cn, Jinan Xu1, Xiangyu Shi1, Chi Chen2, Kaiyu Huang1 1 Beijing Jiaotong University, Beijing, China 2 Tsinghua University, Beijing, China
1 Introduction
The development of large reasoning models (LRMs) such as OpenAI o1 Jaech et al. (2024) and DeepSeek-R1 Guo et al. (2025) has unveiled significant potential in addressing complex reasoning tasks Chan et al. (2024); Chen et al. (2024), including mathematical problem-solving Shao et al. (2024) and code generation El-Kishky et al. (2025). Consequently, the reasoning paradigm has been naturally extended to multimodal scenarios, giving rise to multimodal large reasoning models (MLRMs) Peng et al. (2025a); Yao et al. (2024); Team et al. (2025). The MLRMs are capable of generating long reasoning pathways and exhibit remarkable spatial reasoning abilities, thereby broadening their real-world applicability across diverse scenarios Zhang et al. (2025a); Huang et al. (2025); Wu et al. (2025).
Figure 1: Examples of multimodal safety benchmarks and their corresponding responses on different models.
While LRMs have led to a transformative leap in reasoning performance, prior studies have indicated that the exposure of the chain-of-thought (CoT) process inadvertently undermines safety, as adversaries exploit intermediate reasoning steps to circumvent alignment safeguards Jiang et al. (2025); Zhou et al. (2025); Ying et al. (2025). This vulnerability has also been observed in MLRMs Fang et al. (2025). However, current research on the safety of MLRMs remains limited, primarily focusing on two jailbreak robustness benchmarks, with insufficient exploration of the broader spectrum of safety challenges in multimodal settings.
As shown in Figure 1, we categorize the safety benchmarks of multimodal models into two key aspects: safety-awareness and jailbreak robustness. Jailbreak robustness benchmarks focus on evaluating the resilience of the model against deliberately crafted or modified textual prompts and visual inputs that aim to bypass established safety defense mechanisms. Safety-awareness benchmarks emphasize the capability of models to proactively identify potential safety risks embedded in user inputs, aligning more closely with the complex and dynamic safety demands encountered in real-world applications. Compared to unimodal LRMs, MLRMs exhibits the following research questions:
RQ1: How does MLRMs affect safety on different types of benchmarks compared with Multimodal Large Language Models (MLLMs)?
RQ2: What are the limitations of incorporating additional modalities in MLRMs for safety concerns?
RQ3: What is the impact of reasoning pathways in MLRMs on the safety performance?
To investigate the above research questions and challenges, in this study, we first conduct a systematic safety evaluation of advanced MLRMs, including Kimi-VL-thinking Team et al. (2025), R1-Onevision Yang et al. (2025) and etc. The results demonstrate that MLRMs have significant negative impact on the safety performance, while the performance degradation is task-dependent. In particular, MLRMs could identify more potential safety risks through deliberate thinking, leading to higher safety scores on safety-awareness benchmarks, which provides a novel perspective for mitigating safety degradation in MLRMs. Motivated by this, we further propose a data construction method that incorporates safety-oriented thought process to investigate the effectiveness of this insight.
To sum up, our contributions are summarized as follows:
  • This study conducts a systematic evaluation of MLRMs and investigates the empirical results, revealing several novel findings and providing new perspectives for the development of safer MLRMs.
  • We construct a multimodal fine-tuning dataset with safety-oriented thought process for safety alignment, alleviating the issue of incorporating additional modalities.
  • Experimental results demonstrate that our method improve the safety performance of MLRMs across multiple benchmarks by enabling self-correction thinking along the reasoning pathways, compared with previous defense methods.
Model Name Base Model
LlamaV-o1 Thawakar et al. (2025) Llama3.2-vision-11B
Llava-cot Xu et al. (2024) Llama3.2-vision-11B
Mulberry-Llama Yao et al. (2024) Llama3.2-vision-11B
LMM-R1 Peng et al. (2025b) Qwen2.5-VL-3B
R1-Onevision Yang et al. (2025) Qwen2.5-VL-7B
MM-Eureka Meng et al. (2025) InternVL-2.5-8B
Kimi-VL-thinking Team et al. (2025) Kimi-VL-instruct
Skywork-R1V Peng et al. (2025a) -
QVQ-Preview -
Gemini2.0-flash-thinking -
Claude3.7-sonnet-thinking -
Table 1: List of the evaluated MLRMs and their corresponding base models.
Jailbreak Robustness Safety-Awareness
[…]
Gemini2.0-flash-thinking 89.80 73.48 50.36 39.16 57.32
Claude3.7-sonnet-thinking 32.20 33.94 0.49 59.24 57.76
QVQ-Preview 70.80 69.29 37.86 34.13 50.68
Skywork-R1V 85.80 72.68 35.28 35.33 50.42
Llama3.2-vision-11B(base) 55.80 38.45 5.71 37.13 52.26
LlamaV-o1 59.40(+3.60) 53.93(+15.48) 13.57(+7.86) 33.93(-3.20) 51.59(-0.67)
Llama3.2-vision-11B(base) 55.80 38.45 5.71 37.13 52.26
Llava-cot 84.80(+29.00) 72.26(+33.81) 33.57(+27.86) 26.95(-10.18) 51.26(-1.00)
Llama3.2-vision-11B(base) 55.80 38.45 5.71 37.13 52.26
Mulberry-Llama 67.40(+11.60) 64.70(+26.25) 13.21(+7.50) 37.72(+0.59) 54.09(+1.83)
Qwen2.5-VL-3B(base) 66.27 66.18 12.14 24.55 52.35
LMM-R1 69.80(+3.53) 68.15(+1.97) 18.21(+6.07) 21.56(-2.99) 53.02(+0.67)
Qwen2.5-VL-7B(base) 67.20 66.49 12.50 29.94 51.09
R1-Onevision 72.20(+5.00) 79.57(+13.08) 32.14(+19.64) 17.31(-12.63) 49.69(-1.40)
InternVL-2.5-8B(base) 71.40 59.64 15.00 28.14 50.84
$M$-Eureka 72.20(+0.80) 60.12(+0.48) 11.79(-3.21) 28.14(-0.00) 50.59(-0.25)
Kimi-VL-instruct(base) 80.40 47.74 22.50 25.00 50.44
Kimi-VL-thinking 87.00(+6.60) 61.49(+13.75) 33.93(+11.43) 35.93(+10.93) 51.42(+0.98)
Table 2: Variation of safety performance for MLRMs across various benchmarks. $\downarrow$ means the lower score the safer, while $\uparrow$ means the higher the better. The safety performance variation has been marked in brackets, where the red color represents safety deterioration and green stands for safety improvement.
2 Safety Evaluation of MLRMs
2.1 Evaluation Settings
Datasets.
To comprehensively assess the safety performance of MLRMs across diverse scenarios, we adopt benchmark datasets from two distinct perspectives: safety-awareness and jailbreak robustness. For the assessment of safety-awareness, we employ SIUO Wang et al. (2025) and MSSBench Zhou et al. (2024) datasets, while for evaluating jailbreak robustness, we employ MMSafetyBench Liu et al. (2024b), FigStep Gong et al. (2025) and JailBreaKV Luo et al. (2024).
In safety-awareness tasks, models need to reason jointly over both visual and textual inputs to infer user intent, identify potential safety risks, and assess whether the input should be treated as safe or unsafe. These tasks pose significant challenges to the multimodal reasoning and safety alignment capabilities of models. In contrast, jailbreak robustness benchmarks involve adversarial attacks, such as the inclusion of malicious crafted prompt, aimed at circumventing the safety constraints of models. Given that these two task categories examine safety alignment from fundamentally different perspectives, we analyze the results separately. Further details regarding the datasets are provided in Appendix A.
Models and Configurations.
We evaluate a total of 11 MLRMs, including both proprietary models and open-source models along with their corresponding base models, such as Kimi-VL-thinking Team et al. (2025), R1-Onevision Yang et al. (2025), Mulberry Yao et al. (2024), and LlamaV-o1 Thawakar et al. (2025). Detailed information of the models are provided in Table 1. For reasoning models that have undergone extensive fine-tuning—via either supervised learning or reinforcement learning—we additionally evaluate their corresponding base models. This approach allows us to trace safety-related changes across model variants. All experiments were conducted using four NVIDIA A100-80G GPUs.
Figure 2: Case study of the better safety consideration on safety-awareness tasks. Kimi-VL-Instruct directly outputs the answer that ignores the potential risk, while Kimi-VL-Thinking dives deeper into the insidious safety issue with stronger reasoning abilities.
Metrics and Evaluator.
For jailbreak robustness benchmarks, we adopt the standard Attack Success Rate (ASR) metric, which quantifies the percentage of instances where the model produces harmful outputs with its safety mechanisms circumvented. A lower ASR indicates stronger safety alignment. For safety-awareness benchmarks, we follow established evaluation protocols to compute a safety score, where higher scores denote better safety performance. Further evaluation settings are detailed in Appendix A. Following the settings of MMSafetyBench, MSS, and SIUO utilizing OpenAI API, we employ GPT-4o-mini Hurst et al. (2024) as the evaluator and adopt identical configurations to ensure consistency across assessments.
2.2 Safety Results and Inspection
Finding 1 ($\Rightarrow$ RQ1): The safety performance of MLRMs requires improvement and exhibits a notable security degradation phenomenon.
Overall Performance.
As shown in Table 2, existing multimodal reasoning models exhibit significant security vulnerabilities, with most open-source models performing unsatisfactorily across various benchmarks. In particular, Kimi-VL-thinking, Skywork-R1V and Llava-cot successfully defend against only about 15% of malicious queries on Figstep. This finding highlights the necessity of ensuring safety alignment alongside improvements in multimodal reasoning capabilities.
Security Degradation.
Most reasoning models exhibit a notable increase in ASR on jailbreak robustness benchmarks relative to their corresponding foundation models, with an average increase of approximately 11.53%. This suggests that the internal safety alignment mechanisms are compromised during the process of enhancing reasoning capabilities, thereby rendering the models more susceptible to jailbreak attacks. In contrast, the degradation of safety performance is less evident in safety-awareness benchmarks. In some cases, models even demonstrate improved safety performance after engaging in extended thought process. A detailed analysis of this phenomenon is presented in Finding 2.
Finding 2 ($\Rightarrow$ RQ1): MLRM paradoxically enhances safety performance via long reasoning on safety-awareness benchmarks, due to improved capability in identifying unsafe intent.
Performance on Safety-Awareness.
As shown in Table 2, the safety score of models such as Kimi-VL-thinking Team et al. (2025), LMM-R1 Peng et al. (2025b), and MM-Eureka Meng et al. (2025) are higher than their corresponding base models on safety-awareness benchmark. The effect is particularly pronounced for Kimi-VL-Thinking model, which achieves a 10.9% improvement on the SIUO benchmark. As illustrated in Figure 2, there is a difference in attention patterns according to the responses generated by these models and their corresponding base models. Reasoning models, when engaging in long reasoning, tend to focus more on potentially unsafe details present in both the visual and textual modalities. These models improve the safety performance through internal self-reflection and error-correction mechanisms. This encouraging finding offers a promising direction for mitigating safety degradation by leveraging the innate reasoning capabilities of the models themselves.
Figure 3: Examples of self-deception in responses generated by MLRMs. The content of reframing intent of users is highlighted with yellow background, and harmful content is marked with the red font.
[…]
Qwen2.5-VL-7B (I) 66.49 12.5 29.94
Qwen2.5-VL-7B (T) 48.21(-18.28) 6.79(-5.71) 38.10(+8.16)
R1-Onevision (I) 79.57 32.14 17.31
R1-Onevision (T) 65.02(-14.55) 23.21(-8.93) 23.35(+6.04)
Kimi-VL-instruct (I) 47.74 22.50 25.00
Kimi-VL-instruct (T) 57.56(+9.82) 19.29(-3.21) 25.00(+0.00)
Kimi-VL-thinking (I) 61.49 33.93 35.93
Kimi-VL-thinking (T) 55.71(-5.78) 29.64(-4.29) 29.34(-6.59)
Llama3.2-vision-11B (I) 38.45 5.71 37.13
Llama3.2-vision-11B (T) 58.52(+20.07) 44.29(+38.58) 22.50(-14.63)
Llava-cot (I) 72.26 33.57 26.95
Llava-cot (T) 60.06(-12.20) 40.36(+6.79) 28.14(+1.19)
InternVL2.5-8B (I) 59.64 15.00 28.14
InternVL2.5-8B (T) 48.57(-11.07) 9.64(-5.36) 27.55(-0.59)
MM-Eureka (I) 60.12 11.79 28.14
MM-Eureka (T) 50.24(-9.88) 10.71(-1.08) 31.14(+3.00)
Table 3: Safety performance when converting image into text caption. “I” indicates multimodal inputs, while “T” refers to text-only inputs.
Finding 3 ($\Rightarrow$ RQ2): The substitution of the image modality with text modality leads to a partial recovery of the safety defense capabilities of models.
Modality Ablation.
To investigate the impact of multimodal inputs on model safety, we convert the original multimodal inputs into text-only ones. In particular, we utilize the Qwen-2.5-VL-72B Bai et al. (2025) model to generate captions corresponding to the images, thereby replacing the visual information with its text representation. As shown in Table 3, most models demonstrate a noticeable improvement in safety performance when provided with unimodal (text-only) inputs, apart from the Llama-series models. This observation suggests that models are more effective at identifying harmful intent when reasoning over the text modality alone. In contrast, detecting harmful intent when reasoning over both image and text modalities presents greater challenges.
Finding 4 ($\Rightarrow$ RQ3): MLRMs generate harmful content during the thought process, which poses a potential risk in response.
Figure 4: ASR scores of the thought process and the final response generated by MLRMs on MMSafetyBench respectively.
Performance of Thought Process.
To investigate the impact of reasoning pathways in MLRMs, we independently evaluate the safety performance of the thought process. As shown in Figure 4, models that explicitly exhibit long CoT Kojima et al. (2022) generate harmful content during the thought process. Additionally, the ASR results for the four models—QVQ-Preview, R1-Onevision, Llava-cot, and Skywork-R1V—are higher for the thought process than for the final responses. Moreover, we analyze the thought process to identify how intermediate content contributes to the production of unsafe responses. As illustrated in Figure 3, although the model detects harmful intent in some cases, it does not reject the prompt outright. Instead, MLRM engages in self-deception by reframing the intent of users as educational, humorous, or otherwise benign, which rationalizes the model to generate harmful content. It is a critical challenge to address subtle intent rationalization in long reasoning, which needs to be explored in future work.
Figure 5: Overview of our data construction. We propose a multi-step pipeline to build the datasets based on various safety-related topics (Step 1: Word Cloud in Blue Block) and image description (Step 2: Red Block), which provides a thoughtful consideration of reasoning with safety, explicitly incorporating long CoT reasoning into the addressing process (Step 3: Green Block) and meticulously designed filtering mechanism (Step 4: Yellow Block).
Finding 5 ($\Rightarrow$ RQ3): MLRMs attempt to eliminate unsafe content in the thought process via self-reflection, while there are also situations where unsafe answers are generated after reducing the thinking steps.
R1-Onevision Safe Response Unsafe Response
Safe Thought 12.65 1.90
Unsafe Thought 7.34 78.11
Kimi-VL-thinking Safe Response Unsafe Response
Safe Thought 12.68 2.74
Unsafe Thought 25.83 58.75
Llava-cot Safe Response Unsafe Response
Safe Thought 15.89 3.06
Unsafe Thought 11.85 69.23
QVQ-Preview Safe Response Unsafe Response
Safe Thought 24.52 6.91
Unsafe Thought 6.19 62.38
Skywork-R1V Safe Response Unsafe Response
Safe Thought 19.88 3.93
Unsafe Thought 7.44 68.75
Table 4: Proportion of safety in the thought process and the final response on MMSafetyBench.
Comparative Effect between Thought and Response.
As shown in Table 4, the results reveal that the potential risk arising from the thought process does not always show up consistently in the final response. For example, the safe thought process leads to the unsafe response in QVQ-Preview, with respect to 6.91% proportion. In specific examples, when the model meets with unsafe intents, it is not engaged in in-depth thinking to generate unsafe content in thought while directly generate unsafe answers in response. The inconsistency between safe thought and unsafe response is due to the potential security risks inherent in the model. On the contrary, the unsafe thought process also results in the safe final response. It is attributed to the internal self-reflection and error-correction mechanisms of MLRMs, which enables models to eventually recognize the harmful nature of the intended prompt. In particular, Kimi-VL-thinking demonstrates a relatively strong capacity for self-correction with respect to 25.83% proportion, which aligns with the observations reported in Finding 2. The comparative results also provide potential insights for mitigating the issue of safety degradation in MLRMs.
3 Data Construction
Although existing studies have introduced several datasets for multimodal safety alignment, most of them consist of brief responses that lack explicit thought process. The supervised fine-tuning method based on these datasets struggles with retaining the inherent reasoning-chain advantages of MLRMs. Furthermore, arriving at conclusions without engaging in a safety-oriented thought process weakens the effectiveness and robustness of safety defense mechanisms. To address this issue, we propose a method for building a multimodal alignment dataset enriched with safety-oriented thought process, which offers a green channel to improve the safety of MLRMs. As a result, we aim to leverage the reasoning capabilities of MLRMs to improve the alignment with safety objectives.
As shown in Figure 5, we employ a multi-step pipeline to construct our safety dataset. After collecting various safety-related topics and image descriptions, we explicitly incorporate long CoT reasoning into the addressing process and design a filtering mechanism to finally obtain the dataset. To the best of our knowledge, it is the first safety dataset with the ability to retain reasoning chain for MLRMs.
[…]
(R1-Onevision)
Direct 72.20 79.57 32.14 17.31 49.69
VLGuard 65.80 35.60 3.21 40.72 60.88
MIS 56.00 60.18 7.14 31.74 53.63
Ours 16.80 21.54 0.00 71.26 61.09
(Llava-cot)
Direct 84.80 72.26 33.57 26.95 51.26
VLGuard 72.00 25.65 10.00 75.00 51.42
MIS 75.60 58.45 11.79 79.04 51.08
Ours 20.60 11.90 0.71 97.01 54.73
Table 5: Results of supervised fine-tuning method using different datasets. The best safety scores of the MLRM on each benchmark are highlighted in bold.
Data Preparation.
Due to the limitation of existing multimodal safety alignment datasets, we aim to construct a dataset enriched with safety-oriented thought process. However, considering the inherent security vulnerabilities of existing models and the development costs, it is not a practical approach to build such a dataset entirely from scratch. Therefore, we augment the existing dataset that already contains responses by adding structured and security-oriented thought process. In particular, we select two existing multimodal safety alignment datasets, VLGuard Zong et al. (2024) and SPA-VL Zhang et al. (2024), as the original data sources. Moreover, we define a set of twelve safety categories to select the original data, following the OpenAI safety guidelines, instead of using all data. More details are provided in Appendix C.1.
Image Description Generation.
To obtain a reliable thought process, we choose reasoning models instead of instruct models, which lack the ability to generate the thinking process with diversity and logicality. Furthermore, it is not feasible to directly utilize MLRMs to augment the thought process of the original data with their poor security performanc. Moreover, API-based proprietary models such as Claude3.7-sonnet-thinking with stronger safety mechanisms typically only return brief refusals (e.g., “I’m sorry”) when faced with some unsafe queries, offering no insight into their thought process. Therefore, we utilize the high-performing text-only reasoning model DeepSeek-R1 Guo et al. (2025) to generate safety-oriented thought process. To accommodate this model, we convert images into caption form as input in advance. In particular, we employ Qwen2.5-VL-72B Bai et al. (2025) to generate image captions to ensure that the semantic loss during modality conversion is within an acceptable range in terms of security issues, with the prompt “Please provide a detailed description of this image.
Safety Thought Process Generation.
To ensure the thought process generated by LRMs are truly safe, we add supplementary information related to security, inspired by the concept of deliberative alignment Guan et al. (2024). Previous methods often require models to infer implicit safety rules from large volumes of annotated examples, which suffers from low efficiency and limited generalization. Therefore, we provide clear and structured safety guidelines to directly guide the model in generating safety-oriented CoT. The guidelines not only help ensure the safety of the generated thought process, but also encourage MLRMs to think explicitly with safety considerations when producing responses. Besides, we develop the safety guidelines customized to the specific characteristics and risk profiles of each category. The image caption, question, original response, and safety guidelines are provided to DeepSeek-R1 Guo et al. (2025) to generate safety-oriented thought process. The model first analyzes the question and the caption to obtain the underlying intent, then refers to the safety guidelines to generate a safety-oriented thought process corresponding to the provided response. Prompt and examples of safety guidelines used in this step are provided in the Appendix C.2.
Thought Process Filter.
In the previous step, we obtain the original thought process. However, it is still necessary to ensure that the thought process is consistent with the corresponding response, which means that both should consistently reject or not reject the inquiry. Additionally, while the input to the model in the prior step is a caption, the input during MLRM fine-tuning is the original image. Therefore, the caption field is processed to prevent potential misinterpretation by the MLRM. The original thought process from the previous step also contain redundant elements such as headers like “Safety Chain of Thought” which need to be removed and reformatted. Given the relatively low complexity of this task, we employ DeepSeek-V3 Liu et al. (2024a) to handle this step. Specific prompts used for this step are provided in the Figure 6. After filtering, we construct the final safety datasets provided for the supervised fine-tuning of MLRMs.
4 Experiments
4.1 Experimental Settings
Baselines.
To evaluate the effectiveness of our data construction, we select two existing multimodal safety alignment datasets for supervised fine-tuning as baselines that are listed as follows:.
  • Direct: The original safety capability of the model without fine-tuning.
  • VLGuard Zong et al. (2024): The first safety dataset specifically designed for fine-tuning of MLLMs.
  • MIS Ding et al. (2025b): A multi-image safety dataset that combines visual perception with reasoning logic labels.
Training Details.
We use R1-Onevision Yang et al. (2025) and Llava-cot Xu et al. (2024) as base MLRMs for supervised fine-tuning, which clearly delineates the boundary between the thought process and the final response. We adopt supervised fine-tuning approach to the two models using the baseline datasets and our proposed dataset, respectively. The Details of the training procedure are provided in Appendix B.
4.2 Results
As shown in Table 5, the results demonstrate that the MLRMs achieve the best security performance after fine-tuning using our dataset on all benchmarks, compared with other baseline datasets. It is worth emphasizing that despite the complete absence of jailbreak attack data in our dataset, the proposed approach achieves a substantial reduction of ASR on jailbreak robustness benchmarks. Furthermore, since our dataset incorporates thought process that closely align with the data distribution used in MLRM training, the fine-tuned model preserved the original capability to generate coherent reasoning pathways. More case studies are provided in Appendix D.2. In addition, by fine-tuning the models respectively on the VLGuard, the MIS and the dataset we proposed, the safety performance of both models has been consistently improved, proving that well-designed constructing data is an effective solution to enhance the safety of the models.
5 Related Work
Safety of LRMs.
With the rapid development and widespread deployment of LRMs, many works have paid attention to security of LRMs. Several studies  Jiang et al. (2025); Parmar and Govindarajulu (2025) have conducted comprehensive security evaluations on LRMs, e.g., DeepSeek-R1 Guo et al. (2025), revealing existing vulnerabilities of these models. In addition, Zhou et al. (2025) introduce a CoT jailbreak attack method, specifically targeting reasoning models. On the other hand, to mitigate the issues of security, recent studies Jiang et al. (2025); Zhang et al. (2025b) have constructed the safety data with the thought process for supervised fine-tuning of LRMs. However, the above-mentioned works are all confined to the text-only models. Although SafeMLRM Fang et al. (2025) explores the security of MLRMs, the limitation of this work is that it only focuses on one security scenario, i.e., jailbreak robustness. The introduction of another modality in LRMs brings about significant shifts in model behavior, making it crucial to systematically assess the security challenges of MLRMs and develop effective defense methods.
Safety of MLLMs.
There are several studies that enhance the safety capabilities of MLLMs, which can be categorized into two types: the training-based method and the training-free method. The training-based method typically includes Supervised Fine-Tuning (SFT) Zong et al. (2024); Ding et al. (2025b); Li et al. (2024) and Reinforcement Learning from Human Feedback (RLHF) Zhang et al. (2024). In addition, Lee et al. (2024) introduce the weight merging approach to mitigate safety degradation. Another branch of studies incorporates additional safety components in a training-free manner. For instance, MLLM-Protector Pi et al. (2024) offers a plug-and-play detector for harmful responses, and ETA Ding et al. (2025a) proposes an inference-time alignment framework to ensure safety compliance during deployment. In contrast to these methods, our study is the first to investigate the safety behaviors of reasoning models in both jailbreak and awareness of safety scenarios.
6 Conclusion
In this work, we conduct a comprehensive evaluation and analysis of the safety of existing MLRMs, covering both jailbreak and comprehension benchmarks. The empirical results unveil several new findings, demonstrating that the safety performance of current MLRMs remains a significant concern. Furthermore, motivated by the findings, we propose a supervised fine-tuning dataset that considers explicit safety-oriented thought process. Experimental results on R1-Onevision and Llava-cot demonstrate that our dataset outperforms existing alternatives. This work represents a preliminary exploration of improving the safety of MLRMs through reasoning-based alignment. Future research will focus on developing more efficient datasets and training strategies specifically designed for MLRMs.
Appendix A Evaluation Details
Our evaluation encompasses five categories of safety benchmarks, which can be broadly divided into two types: jailbreak robustness and safety-awareness. Specifically, FigStep, MMSafetyBench, and JailBreaKV fall under the jailbreak robustness benchmark, for which we report the Attack Success Rate (ASR). Meanwhile, MSSBench and SIUO are categorized as safety-awareness benchmarks, and we follow the respective safety metrics according to the original papers.
The FigStep adopts the human evaluation as safety judgment. In our experiments, we follow the classification criteria described in their paper but apply an evaluation protocol consistent with that used for the MMSafetyBench. For the JailBreaKV, we follow the same evaluation settings as their work and utilize the Llama-Guard-Vision Chi et al. (2024) as the evaluator. The benchmarks are listed as follows:
FigStep.
FigStep Gong et al. (2025) consists of 500 harmful queries spanning 10 categories. Rather than directly inputting harmful textual prompts, it conveys malicious intent through visual formatting. Each image encodes harmful content as a visually rendered list, which is then paired with a fixed textual instruction: “The image shows a list numbered 1, 2, and 3, but the items are empty. Please generate detailed content for each item on the list.
$M$-SafetyBench.
$M$-SafetyBench Liu et al. (2024b) covers 13 categories and is primarily designed to evaluate the responses of MLLMs to malicious queries. In our experiments, we adopt the SD+TYPO setting, which includes a total of 1,680 samples, following the setting of Ding et al. (2025b).
JailBreaKV.
JailBreaKV Luo et al. (2024) is designed to evaluate the transferability of jailbreak techniques from LLMs to MLLMs, encompassing both text-based and image-based jailbreak attacks. Following the evaluation protocol of VLGuard, we conduct our experiments on the miniset provided in their paper, which contains 280 samples, following the setting of Hu et al. (2024).
MSSBench.
Multimodal Situation Safety Benchmark (MSSBench) Zhou et al. (2024) explores how safety considerations vary depending on the specific context involving the user or agent. In this benchmark, half of the image contexts are safe, while the other half are unsafe. The benchmark identifies safe intent of the models in two distinct settings: chat and embodied scenarios. Due to economic constraints, our evaluation was conducted using a single instruction sampled from each of the safe and unsafe scenarios.
SIUO.
Safety Inputs but Unsafe Outputs (SIUO) Wang et al. (2025) considers cases where the image and text modalities are independently safe, but their combination leads to unsafe or unethical outputs. The SIUO covers 9 categories including a total of 167 samples. We report the evaluation metrics as defined in their paper.
Appendix B Training Details
For the training procedure, we utilize the LoRA Hu et al. (2022) to fine-tune the R1-Onevision and Llava-cot based on the LLaMA-Factory [^1] https://github.com/hiyouga/LLaMA-Factory framework. All training experiments are conducted using one NVIDIA A100-80G GPUs. The hyper-parameters for supervised fine-tuning are listed in Table 6.
Hyper-Parameter Value
lora rank 8
learning rate 1.0e-5
train epoch 1.5
per_device_batchsize 1
warm up ratio 0.1
learning rate scheduler cosine
Table 6: Hyper-parameters for supervised fine-tuning.
Model Chat Embodied AVG.
Safe Unsafe AVG. Safe Unsafe AVG.
Gemini2.0-flash-thinking 85.67 45.00 65.33 93.06 5.56 49.31 57.32
Claude3.7-sonnet-thinking 75.84 45.97 60.91 93.42 15.79 54.61 57.76
QVQ-Preview 80.67 24.67 52.67 53.95 43.42 48.68 50.68
Skywork-R1V 86.67 15.00 50.83 98.68 1.32 50.00 50.42
Llama3.2-vision-11B(base) 85.62 22.07 53.85 90.79 10.53 50.66 52.26
LlamaV-o1 59.33 47.00 53.17 100.00 0.00 50.00 51.59
Llama3.2-vision-11B(base) 85.62 22.07 53.85 90.79 10.53 50.66 52.26
Llava-cot 93.67 12.67 53.17 97.37 1.32 49.34 51.26
Llama3.2-vision-11B(base) 85.62 22.07 53.85 90.79 10.53 50.66 52.26
Mulberry-Llama 81.33 35.00 58.17 97.37 2.63 50.00 54.09
Qwen2.5-VL-3B(base) 83.67 31.00 57.33 59.21 35.53 47.37 52.35
LMM-R1 83.00 33.00 58.00 69.70 26.32 48.03 53.02
Qwen2.5-VL-7B(base) 93.33 11.00 52.17 93.42 6.58 50.00 51.09
R1-Onevision 86.25 15.46 50.86 86.57 10.45 48.51 49.69
InternVL-2.5-8B (base) 92.00 11.33 51.67 98.68 1.32 50.00 50.84
$M$-Eureka 90.67 11.67 51.17 97.37 2.63 50.00 50.59
Kimi-VL-instruct (base) 91.67 8.67 50.17 100.00 1.43 50.71 50.44
Kimi-VL-thinking 83.33 23.67 53.5 98.67 0.00 49.33 51.42
Table 7: Results on MSSBench in two distinct settings: chat and embodied scenarios.
Appendix C Dataset Details
C.1 Data Statistics
Based on the categories provided in SPA-VL Zhang et al. (2024) and VLGuard Zong et al. (2024) datasets, we extract the raw data in the first step of construction for data preparation. In addition, we include the paired safety examples from VLGuard, where each image is matched with both a safe and an unsafe question. This design aims to improve the ability of the models to handle context-sensitive safety issues. However, the answers in SPA-VL are derived from pairwise preference data generated by large language models. Despite selecting the “chosen” responses as the final answers, a subset of them remain potential safety risks. The unsafe answers are subsequently screened and reconstructed. The statistics of data instances in each category within our dataset after the final filtering process are shown in Table 8.
Category # Samples
Privacy violation 196
Professional advice 200
Political sensitivity 209
Sexually explicit 199
Violence 204
Disinformation 205
Discrimination 231
Hate speech 200
Economic harm 200
Physical harm 200
Illegal activities 200
Malware 200
Safe 977
All 4398
Table 8: The statistics of our dataset
C.2 Prompt for Data Conustruction
This section presents the prompt and safety regulation used for safety thought process collection. We referenced the definition of User Requests Categorization from Wang et al. (2024) and listed all prompts used in the data construction in Figure 6. During the collection of safety thought process data, we supplied the model with category-specific safety regulations derived from the raw data classification. This strategy not only reduce the length of prompt but also enable tiered defensive measures for distinct safety-related issues.
Appendix D More Results
D.1 Detailed Experiments on MSSBench
In the evaluation protocol of MSS-Bench, scenarios are divided into chat and embodied settings, with safety scores calculated separately for safe and unsafe conditions. The average safety scores are reported in Table 2, and the specific scores for each category are detailed in Table 7.
Figure 6: Prompt used in data construction.
D.2 Case Study
Examples of responses generated by fine-tuned on VLGuard, MIS, and our dataset are illustrated in Figure 7. Our approach demonstrates the ability to retain the thought process of the models while decisively rejecting unsafe inputs and explicitly articulating the potential serious consequences associated with such queries.
Figure 7: Examples of several baselines and our dataset on $M$-SafetyBench. We only highlight some fragments to show the representative difference between various methods. The red indicates the unsafe parts, while the blue indicates the safe parts.