### v4

# Think in Safety: Unveiling and Mitigating Safety Alignment Collapse in Multimodal Large Reasoning Model

Abstract
The rapid development of Multimodal Large Reasoning Models (MLRMs) has demonstrated broad application potential, yet their safety and reliability remain critical concerns that require systematic exploration. To address this gap, we conduct a comprehensive and systematic safety evaluation of 13 MLRMs across 5 benchmarks and unveil prevalent safety degradation phenomena in most advanced models. Moreover, our analysis reveals distinct safety patterns across different benchmarks: significant safety degradation is observed across jailbreak robustness benchmarks, whereas safety-awareness benchmarks demonstrate less pronounced degradation. In particular, the long thought process in some scenarios even enhances safety performance. Therefore, it is a potential approach to address safety issues in MLRMs by leveraging the intrinsic reasoning capabilities of the model to detect unsafe intent. To operationalize this insight, we construct a multimodal tuning dataset that incorporates a safety-oriented thought process. Experimental results from fine-tuning existing MLRMs with this dataset effectively enhance the safety on both jailbreak robustness and safety-awareness benchmarks. This study provides a new perspective for developing safe MLRMs. [^1] Our dataset is available at https://github.com/xinyuelou/Think-in-Safety.
Warning: this paper contains example data that may be offensive or harmful.
Think in Safety: Unveiling and Mitigating Safety Alignment Collapse in Multimodal Large Reasoning Model
Xinyue Lou1,2, You Li1,2, Jinan Xu1,2, Xiangyu Shi1,2, Chi Chen3, Kaiyu Huang1,2222Kaiyu Huang is the corresponding author. 1Key Laboratory of Big Data & Artificial Intelligence in Transportation (Beijing Jiaotong University), Ministry of Education 2School of Computer Science and Technology, Beijing Jiaotong University 3Tsinghua University {louxinyue,kyhuang}@bjtu.edu.cn
1 Introduction
The development of Large Reasoning Mode-ls (LRMs) such as OpenAI o1 Jaech et al. (2024) and DeepSeek-R1 Guo et al. (2025) has unveiled significant potential in addressing complex reasoning tasks Chan et al. (2024); Chen et al. (2024), including mathematical problem-solving Shao et al. (2024) and code generation El-Kishky et al. (2025). Consequently, the reasoning paradigm has been naturally extended to multimodal scenarios, giving rise to Multimodal Large Reasoning Models (MLRMs) Peng et al. (2025a); Yao et al. (2024); Team et al. (2025). The MLRMs are capable of generating long reasoning pathways and exhibit remarkable spatial reasoning abilities, thereby broadening their real-world applicability across diverse scenarios Zhang et al. (2025a); Huang et al. (2025); Wu et al. (2025).
Figure 1: Examples of multimodal safety benchmarks and their corresponding responses on different models.
While LRMs have led to a transformative leap in reasoning performance, prior studies have indicated that the exposure of the Chain-of-Thought (CoT) process inadvertently undermines safety, as adversaries exploit intermediate reasoning steps to circumvent alignment safeguards Jiang et al. (2025); Zhou et al. (2025); Ying et al. (2025). This vulnerability has also been observed in MLRMs Fang et al. (2025). However, current research on the safety of MLRMs remains limited, primarily focusing on jailbreak robustness benchmarks Gong et al. (2025); Liu et al. (2024b), with insufficient exploration of the broader spectrum of safety challenges in multimodal settings.
As shown in Figure 1, we categorize the safety benchmarks of multimodal models into two key aspects: safety-awareness and jailbreak robustness. Jailbreak robustness benchmarks Gong et al. (2025); Liu et al. (2024b); Luo et al. (2024) focus on evaluating the resilience of the model against deliberately crafted or modified textual prompts and visual inputs that aim to bypass established safety defense mechanisms. Safety-awareness benchmarks Wang et al. (2025); Zhou et al. (2024) emphasize the capability of models to proactively identify potential safety risks embedded in user inputs, aligning more closely with the complex and dynamic safety demands encountered in real-world applications. Compared to unimodal LRMs, MLRMs exhibits the following research questions:
RQ1: How do MLRMs affect safety on different types of benchmarks compared with Multimodal Large Language Models (MLLMs)?
RQ2: What are the risks of incorporating additional modalities in MLRMs for safety concerns?
RQ3: What is the impact of reasoning pathways in MLRMs on the safety performance?
To investigate the above research questions and challenges, in this study, we first conduct a systematic safety evaluation of advanced MLRMs, including Kimi-VL-Thinking Team et al. (2025), R1-Onevision Yang et al. (2025) , etc. The results demonstrate that MLRMs have significant negative impact on the safety performance, while the performance degradation is task-dependent. Furthermore, we found that MLRMs could identify more potential safety risks through deliberate thinking, leading to higher safety scores on safety-awareness benchmarks, which provides a novel perspective for mitigating safety degradation in MLRMs. Motivated by this, we further propose a data construction method that incorporates safety-oriented thought process to investigate the effectiveness of this insight.
To sum up, our contributions are summarized as follows:
  • This study conducts a systematic safety evaluation of MLRMs and investigates the empirical results, revealing several novel findings and providing new perspectives for the development of safer MLRMs.
  • We construct a multimodal fine-tuning dataset with safety-oriented thought process for safety alignment, alleviating the issue associated with incorporating additional modalities.
  • Experimental results demonstrate that our method improves the safety performance of MLRMs across multiple benchmarks by enabling self-correction thinking along the reasoning pathways, compared with previous defense methods.
  Jailbreak Robustness Safety-Awareness
[…]
Table 1: Variation of safety performance for MLRMs across various benchmarks. $\downarrow$ means the lower score the safer, while $\uparrow$ means the higher the better. The safety performance variation has been marked in brackets, where the red color represents safety deterioration and green stands for safety improvement.
2 Safety Evaluation of MLRMs
2.1 Evaluation Settings
Datasets.
To comprehensively assess the safety performance of MLRMs across diverse scenarios, we adopt benchmark datasets from two distinct perspectives: safety-awareness and jailbreak robustness. For assessment of safety-awareness, we employ SIUO Wang et al. (2025) and MSSBench Zhou et al. (2024) datasets, while for evaluating jailbreak robustness, we employ MM-SafetyBench Liu et al. (2024b), FigStep Gong et al. (2025) and JailBreaKV Luo et al. (2024).
In safety-awareness tasks, models need to jointly reason over both visual and textual inputs to infer user intent, identify potential safety risks, and assess whether the input should be treated as safe or unsafe. These tasks pose significant challenges to the multimodal reasoning and safety alignment capabilities of models. In contrast, jailbreak robustness benchmarks involve adversarial attacks, such as the inclusion of maliciously crafted prompt, aimed at circumventing the safety constraints of models. Given that these two task categories examine safety alignment from different perspectives, we analyze these results separately. Further details regarding the datasets are provided in Appendix A.1.
Models and Configurations.
We evaluate a total of 13 MLRMs, including both proprietary models and open-source models along with their corresponding base models, such as Kimi-VL-Thinking Team et al. (2025), R1-Onevision Yang et al. (2025), Mulberry Yao et al. (2024), and LlamaV-o1 Thawakar et al. (2025). Detailed information of the models are provided in Appendix A.3. For reasoning models that have undergone extensive fine-tuning through either supervised learning or reinforcement learning, we additionally evaluate their corresponding base models. This approach allows us to trace safety-related changes across model variants. All experiments are conducted using two NVIDIA A100-80G GPUs.
Metrics and Evaluator.
For jailbreak robustness benchmarks, we adopt the standard Attack Success Rate (ASR) metric, which quantifies the percentage of instances where the model produces harmful outputs with its safety mechanisms circumvented. A lower ASR indicates stronger safety alignment. For safety-awareness benchmarks, we follow established evaluation protocols to compute a safety score, where higher scores denote better safety performance. Further evaluation settings are detailed in Appendix A.1. Following the settings of MM-SafetyBench Liu et al. (2024b), MSSBench Zhou et al. (2024), and SIUO Wang et al. (2025) utilizing OpenAI API, we employ GPT-4o-mini Hurst et al. (2024) as the evaluator and adopt identical configurations to ensure consistency across assessments.
Figure 2: Case study of the better safety consideration on safety-awareness tasks. Kimi-VL directly outputs the answer that ignores the potential risk, while Kimi-VL-Thinking dives deeper into the insidious safety issue with stronger reasoning abilities. The red indicates the unsafe parts, while the green indicates the content related to potential risks identified by reasoning models.
2.2 Safety Results and Inspection
Finding 1 ($\Rightarrow$ RQ1): The safety performance of MLRMs requires improvement and exhibits a notable safety degradation phenomenon.
Overall Performance.
As shown in Table 1, existing multimodal reasoning models exhibit significant safety vulnerabilities, with most open-source models performing unsatisfactorily across various benchmarks. In particular, Kimi-VL-Thinking, Skywork-R1V and LLaVA-CoT successfully defend against only about 15% of malicious queries on FigStep. This finding highlights the necessity of ensuring safety alignment alongside improvements in multimodal reasoning capabilities.
Safety Degradation.
Most reasoning models exhibit a notable increase in ASR on jailbreak robustness benchmarks relative to their corresponding foundation models, with an average increase of approximately 10.57%. This suggests that the internal safety alignment mechanisms are compromised during the process of enhancing reasoning capabilities, thereby rendering the models more susceptible to jailbreak attacks. In contrast, the degradation of safety performance is less evident in safety-awareness benchmarks. In some cases, models even demonstrate improved safety performance after engaging in extended thought process. A detailed analysis of this phenomenon is presented in Finding 2.
Finding 2 ($\Rightarrow$ RQ1): MLRMs paradoxically enhances safety performance via long reasoning on safety-awareness benchmarks, due to improved capability in identifying unsafe intent.
[…]
Qwen2.5-VL-7B 66.49 12.5 29.94
     -w/o image 48.21(-18.28) 6.79(-5.71) 38.10(+8.16)
R1-Onevision 79.57 32.14 17.31
     -w/o image 65.02(-14.55) 23.21(-8.93) 23.35(+6.04)
Kimi-VL 47.74 22.50 25.00
     -w/o image 57.56(+9.82) 19.29(-3.21) 25.00(+0.00)
Kimi-VL-Thinking 61.49 33.93 35.93
     -w/o image 55.71(-5.78) 29.64(-4.29) 29.34(-6.59)
Llama 3.2-vision-11B 38.45 5.71 37.13
     -w/o image 58.52(+20.07) 44.29(+38.58) 22.50(-14.63)
LLaVA-CoT 72.26 33.57 26.95
     -w/o image 60.06(-12.20) 40.36(+6.79) 28.14(+1.19)
InternVL2.5-8B 59.64 15.00 28.14
     -w/o image 48.57(-11.07) 9.64(-5.36) 27.55(-0.59)
$M$-Eureka 60.12 11.79 28.14
     -w/o image 50.24(-9.88) 10.71(-1.08) 31.14(+3.00)
Table 2: Safety performance when converting image into text caption. The safety performance variation has been marked in brackets, where the red color represents safety deterioration and green stands for safety improvement. MMSafe is the abbreviation for $M$-SafetyBench.
Figure 3: ASR scores of the thought process and the final answer generated by MLRMs on FigStep respectively.
Performance on Safety-Awareness.
As shown in Table 1, the safety score of models such as Kimi-VL-Thinking Team et al. (2025), Mixed-R1 Xu et al. (2025), and SophiaVL-R1  Fan et al. (2025) are higher than their corresponding base models on safety-awareness benchmark. The effect is particularly pronounced for SophiaVL-R1 model, which achieves an 11.38% improvement on the SIUO benchmark. As illustrated in Figure 2, there is a difference in focusing center according to the responses generated by these models and their corresponding base models. Reasoning models, when engaging in long reasoning, tend to focus more on potentially unsafe details present in both the visual and textual modalities. These models improve the safety performance through internal self-reflection and error-correction mechanisms. This encouraging finding offers a promising direction for mitigating safety degradation by leveraging the innate reasoning capabilities of the models themselves.
Figure 4: Examples of self-deception in responses generated by MLRMs. The content of reframing intent of users is highlighted with yellow background, and harmful content is marked with the red font.
Finding 3 ($\Rightarrow$ RQ2): The substitution of the image modality with text modality leads to a partial recovery of the safety defense capabilities of models.
Modality Ablation.
To investigate the impact of multimodal inputs on model safety, we convert the original multimodal inputs into text-only ones. In particular, we utilize the Qwen-2.5-VL-72B Bai et al. (2025) model to generate captions corresponding to the images, thereby replacing the visual information with its text representation. As shown in Table 2, most models demonstrate a noticeable improvement in safety performance when provided with unimodal (text-only) inputs, apart from the Llama-series models. This observation suggests that models are more effective at identifying harmful intent when reasoning over the text modality alone. In contrast, detecting harmful intent when reasoning over both image and text modalities presents greater challenges.
  R1-Onevision Safe Answer Unsafe Answer
Safe Thought 13.60% 29.40%
Unsafe Thought 14.20% 42.80%
Kimi-VL-Thinking Safe Answer Unsafe Answer
Safe Thought 8.40% 30.40%
Unsafe Thought 4.60% 48.20%
LLaVA-CoT Safe Answer Unsafe Answer
Safe Thought 8.80% 30.20%
Unsafe Thought 6.40% 54.60%
QVQ-Preview Safe Answer Unsafe Answer
Safe Thought 24.60% 46.60%
Unsafe Thought 4.40% 24.20%
Skywork-R1V Safe Answer Unsafe Answer
Safe Thought 9.40% 37.60%
Unsafe Thought 4.80% 48.20%
Table 3: Proportion of safety in the thought process and the final answer on FigStep.
Finding 4 ($\Rightarrow$ RQ3): MLRMs generate harmful content during the thought process, which poses a potential risk in response.
Performance of Thought Process.
To investigate the impact of reasoning pathways in MLRMs, we independently evaluate the safety performance of the thought process. As shown in Figure 3, the proportion of unsafe thought process generated by the four models (R1-Onevision, LLaVA-CoT, Kimi-VL-Thinking and Skywork-R1V) all exceeds 50 percent. Moreover, we analyze the thought process to identify how intermediate content contributes to the production of unsafe answers. As illustrated in Figure 4, although the model detects harmful intent in some cases, it does not reject the prompt outright. Instead, MLRM engages in self-deception by reframing the intent of users as educational, humorous, or otherwise benign, which rationalizes the model to generate harmful content. It is a critical challenge to address subtle intent rationalization in long reasoning, which needs to be explored in future work.
Figure 5: Examples of the safety inconsistency between the thought process and the final answer. The harmful content is marked with the red font.
Figure 6: Overview of our data construction pipeline. We propose a multi-stage pipeline to build the datasets based on various safety-related topics (Step 1: Word Cloud in Blue Block) and image description (Step 2: Red Block), which provides a thoughtful consideration of reasoning with safety, explicitly incorporating long CoT reasoning into the addressing process (Step 3: Green Block) and meticulously designed filtering mechanism (Step 4: Yellow Block).
Finding 5 ($\Rightarrow$ RQ3): MLRMs attempt to eliminate unsafe content in the thought process via self-reflection, while there are also situations where unsafe answers are generated even if the thought process is safe.
Comparative Effect between Thought and Answer.
As shown in Table 3, the results reveal that the potential risk arising from the thought process does not always show up consistently in the final answer. For example, the safe thought process leads to the unsafe answer in QVQ-Preview, with respect to 46.60% proportion. As shown in Figure 5(a), in specific examples, when the model encounters unsafe intents, the reasoning process focuses merely on the directivity for answering the question, without generating explicitly harmful actions. The inconsistency between safe thought and unsafe answer is due to the potential safety risks inherent in the model. On the contrary, the unsafe thought process also results in the safe final answer. It is attributed to the internal self-reflection and error-correction mechanisms of MLRMs, which enables models to eventually recognize the harmful nature of the intended prompt as shown in Figure 5(b). The comparative results suggesting that we need to not only eliminate unsafe content in the reasoning process or answer section, but also ensure consistency between the thought process and the final answer.
3 Data Construction
Although existing studies have introduced several datasets for multimodal safety alignment, most of them Zong et al. (2024); Ding et al. (2025b) consist of brief responses that lack explicit thought process. The fine-tuning method based on these datasets struggles with retaining the inherent reasoning-chain advantages of MLRMs. Furthermore, arriving at conclusions without engaging in a safety-oriented thought process weakens the effectiveness and robustness of safety defense mechanisms. To address these issues, we propose a method for building a multimodal alignment dataset named TiS (Think in Safety), distinguished by its safety-oriented thought process and various safety-related topics. Through this method, we aim to leverage the innate reasoning capabilities of MLRMs to improve its safety alignment.
As shown in Figure 6, we employ a multi-stage pipeline to construct our safety alignment dataset TiS. We begin by collecting safety-related topics and generating image captions, then explicitly incorporating long CoT reasoning into question answering. After a filtering procedure, we finally obtain the dataset. To the best of our knowledge, TiS is the first alignment dataset with the ability to retain reasoning chain for MLRMs. We provide a detailed illustration of the entire pipeline in Appendix B.1.
  Jailbreak Robustness Safety-Awareness
[…]
Direct 72.20 79.57 32.14 17.31 48.94
VLGuard 36.60(-35.60) 27.21(-52.36) 0.36(-31.78) 45.51(+28.20) 63.86(+14.92)
MIS 68.00(-4.20) 46.02(-33.55) 2.14(-30.00) 41.92(+24.61) 65.61(+15.67)
SPA-VL 30.00(-42.20) 35.30(-44.27) 0.36(-31.78) 42.51(+25.20) 63.60(+14.66)
TiS (Ours) 15.80(-56.40) 21.79(-57.78) 0.00(-32.14) 71.26(+53.95) 66.31(+17.37)
-w/o thought 20.80(-51.40) 16.37(-63.20) 0.00(-32.14) 63.47(+46.14) 63.92(+14.98)
[…]
Direct 84.80 72.26 33.57 26.95 51.67
VLGuard 62.60(-22.20) 14.76(-57.50) 5.36(-28.21) 61.68(+34.73) 53.40(+1.73)
MIS 59.60(-25.20) 45.48(-26.78) 2.86(-3.71) 52.10(+25.15) 54.90(+3.23)
SPA-VL 41.80(-43.00) 43.63(-28.63) 0.71(-32.86) 65.27(+38.32) 57.63(+5.96)
TiS (Ours) 12.20(-72.60) 8.87(-63.39) 0.00(-33.57) 74.85(+47.90) 59.17(+7.50)
-w/o thought 28.60(-56.20) 6.67(-65.59) 2.14(-31.43) 67.66(+40.71) 53.96(+2.09)
Table 4: Results of supervised fine-tuning method using different datasets. The best safety scores of the MLRM on each benchmark are highlighted in bold and the second-best are highlighted in underline.
Data Preparation.
Due to the limitation of existing multimodal safety alignment datasets, we aim to construct a dataset enriched with safety-oriented thought process. However, considering the inherent safety vulnerabilities of existing models and the resource constraints, constructing such a dataset entirely from scratch is largely impractical. Therefore, we augment existing datasets with short responses by adding structured and safety-oriented thought process. Specifically, we select two multimodal safety alignment datasets as our original data sources: VLGuard Zong et al. (2024) and SPA-VL Zhang et al. (2024). Rather than directly utilizing the entire datasets, we retain only a subset of instances to ensure a balanced distribution across diverse safety-related topics.
Image Description Generation.
To obtain a reliable thought process, we utilize the text-only reasoning model Deepseek-R1 Guo et al. (2025). This is because multimodal instruction models often struggle to produce coherent reasoning, while MLRMs tend to suffer from alignment collapse under various conditions. Meanwhile, close-source proprietary models are equipped with defensive safety guardrail, which hinders access to their internal reasoning processes particularly for safety-related questions. To address the modality gap inherent in this text-only models, the visual content is converted into detailed image captions using Qwen2.5-VL-72B Bai et al. (2025).
  R1-Onevision Safe Answer Unsafe Answer
Safe Thought 13.60% 29.40%
Unsafe Thought 14.20% 42.80%
R1-Onevision+TiS Safe Answer Unsafe Answer
Safe Thought 81.60% 15.60%
Unsafe Thought 2.60% 0.20%
LLaVA-CoT Safe Answer Unsafe Answer
Safe Thought 8.80% 30.20%
Unsafe Thought 6.40% 54.60%
LLaVA-CoT+TiS Safe Answer Unsafe Answer
Safe Thought 81.40% 11.80%
Unsafe Thought 6.40% 0.40%
Table 5: Proportion of safety in the thought process and the final answer of fine-tuned model using our dataset on FigStep.
Safety Thought Process Generation.
To ensure the thought process generated by LRMs is aligned with human values, we explicitly add safety guidelines in the generation process, inspired by the concept of deliberative alignment Guan et al. (2024). Previous methods often require models to infer implicit safety rules from large volumes of training examples, which suffers from low efficiency and limited generalization. By contrast, we provide clear and structured safety guidelines to directly guide the model in generating safety-oriented thought process, encouraging MLRMs to think explicitly with safety considerations when producing responses. Besides, we develop the safety guidelines customized to the specific characteristics and risk profiles of each category. More details and prompt used in this step are provided in Appendix B.3.
Thought Process Filtering.
After obtaining the safety-oriented reasoning processes, we perform an additional modification and filtering step to ensure the quality and correctness of the data, making them better suited for safety alignment. This step addresses issues such as inaccuracies (e.g., responses that explicitly state “according to the caption”) and misalignments in reasoning. We collaboratively employ Deepseek-V3 Liu et al. (2024a) alongside human annotators to assess data quality and filter inappropriate instances.
4 Experiments
4.1 Experimental Settings
Baselines.
To compare the effectiveness of our proposed dataset, we select three multimodal safety alignment datasets (MIS Ding et al. (2025b), VLGuard Zong et al. (2024), SPA-VL Zhang et al. (2024)) as our considered baselines. More details are listed in Appendix C.1.
Training Details.
We use R1-Onevision Yang et al. (2025) and LLaVA-CoT Xu et al. (2024) as base MLRMs for safety alignment training, which demonstrates competitive performance in multimodal reasoning tasks yet falls short in safety performance. We conduct comprehensive supervised fine-tuning on both models with all above datasets and our proposed one with detailed reasoning process. The details of the training configuration are provided in Appendix C.2.
4.2 Results
As shown in Table 4, both R1-Onevision and LLaVA-CoT demonstrate improved safety alignment after fine-tuning on TiS, substantially outperforming prior datasets. Specifically, our dataset TiS enhances safety performance on FigStep and SIUO by at least 10% compared to the best alternative baseline, effectively enabling the models to leverage their reasoning capabilities for deeper analysis and unsafe intention detection. Furthermore, since our dataset incorporates thought process that closely align with the data distribution used in MLRM training, the fine-tuned model preserved the original capability to generate coherent reasoning pathways. More case studies are provided in Appendix D.4.
4.3 Discussion
Ablation Studies.
To investigate whether the proposed thought process can further enhance the safety performance of MLRMs, we conduct experiments by removing the thought component from TiS dataset, retaining only the answer portion. As shown in Table 4, retaining the thought process leads to more substantial improvements in safety performance compared to using answer-only data, with the exception of MM-SafetyBench. For certain instances of MM-SafetyBench, the model’s long reasoning leads it to categorize the input as neutral, which leads to the increase of ASR.The answer-only data also exhibits favorable performance, suggesting that the construction of more comprehensive and higher-quality datasets facilitates the safe fine-tuning of MLRMs. However, this can lead to generated responses that not only lack the thought process but also consist solely of brief replies such as “I’m sorry, I can’t assist it.”.More case studies are provided in Appendix D.3. Furthermore, this also demonstrates the feasibility of leveraging the reasoning capabilities of MLRMs to enhance the alignment with safety objectives.
Analysis of Thought Process.
To further assess the impact of Tis on enhancing the safety of internal reasoning, we separately evaluate the safety of both thought process and final answer, as shown in Table 5. The results demonstrate the effectiveness of TiS in safeguarding both thought and answer response, with a substantial reduction in the ASR across all categories. Notably, TiS effectively reduces the proportion of cases in which both thought process and final answer are unsafe, decreasing from 42.80% to 0.20% on R1-Onevision, and from 54.60% to 0.40% on LLaVA-CoT. Furthermore, in scenarios where unsafe content appears solely in the intermediate thinking, our fine-tuning approach also achieves a marked improvement, reducing such instances from 14.20% to 2.60%. These results indicate that TiS largely enhances the reliability and safety of the model’s internal reasoning process.
5 Related Work
Safety of LRMs.
With the rapid development and widespread deployment of LRMs, many works have paid attention to safety of LRMs. Several studies  Jiang et al. (2025); Parmar and Govindarajulu (2025) have conducted comprehensive safety evaluations on LRMs, e.g., DeepSeek-R1 Guo et al. (2025), revealing existing vulnerabilities of these models. In addition, Zhou et al. (2025) introduce a CoT jailbreak attack method, specifically targeting reasoning models. On the other hand, to mitigate the issues of safety, recent studies Jiang et al. (2025); Zhang et al. (2025b) have constructed the safety data with the thought process for supervised fine-tuning of LRMs. However, the above-mentioned works are all confined to the text-only models. Although SafeMLRM Fang et al. (2025) explores the safety of MLRMs and reveals three critical findings such as Reasoning Tax, the limitation of this work is that it only focuses on one safety scenario, i.e., jailbreak robustness. Our study is the first to investigate the safety behaviors of multimodal reasoning models in both jailbreak and awareness of safety scenarios and reveals more findings beyond Reasoning Tax.
Safety of MLLMs.
Current studies enhancing the safety capabilities of MLLMs can be categorized into two types: the training-based method and the training-free method. The training-based method typically includes Supervised Fine-Tuning (SFT) Zong et al. (2024); Ding et al. (2025b); Li et al. (2024) and Reinforcement Learning from Human Feedback (RLHF) Zhang et al. (2024). In addition, Lee et al. (2024) introduce the weight merging approach to mitigate safety degradation. Another branch of studies incorporates additional safety components in a training-free manner. For instance, MLLM-Protector Pi et al. (2024) offers a plug-and-play detector for harmful responses, ECSO Gou et al. (2024) converts the image inputs into text to exploit the safety of pre-aligned LLMs and ETA Ding et al. (2025a) proposes an inference-time alignment framework to ensure safety compliance. In contrast to these methods, our study is the first to propose potential solutions to address the issue of safety degradation of MLRMs.
6 Conclusion
In this work, we systematically evaluate and analyze the safety performance of existing MLRMs, covering both jailbreak robustness and safety-awareness benchmarks. The empirical results unveil several new findings, demonstrating that the safety performance of current MLRMs remains a significant concern. Furthermore, motivated by the findings, we propose a supervised fine-tuning dataset that considers explicit safety-oriented thought process. Experimental results on R1-Onevision and LLaVA-CoT demonstrate that our dataset outperforms existing alternatives. This work represents a preliminary exploration of improving the safety of MLRMs through reasoning-based alignment. Future research will focus on developing more efficient datasets and training strategies specifically designed for MLRMs.
Limitations
Our study primarily employs MLLMs as evaluative judges due to considerations of cost-efficiency and scalability. However, relying solely on MLLMs may compromise the accuracy of safety assessments, particularly in cases involving subtle forms of unsafe content or where the model fails to correctly interpret output response. Additionally, our evaluation includes only a collection of representative MLRMs, which does not capture the full diversity and reasoning capabilities of the broader range of available MLRMs.
Ethical Considerations
In this paper, we primarily focus on investigation on the safety evaluation of MLRMs. All experiments are conducted using publicly released datasets in a controlled setting, thereby avoiding the creation or propagation of new harmful content. We highlight that the goal of our work is to reveal severe safety essue exhibited by MLRMs. Moreover, we designed TiS dataset to support the development of safer MLRMs without raising ethical concerns.
Acknowledgments
The research work descried in this paper has been supported by the National Nature Science Foundation of China (No. 62376019, 62476023, 61976015, 61976016, 61876198 and 61370130), and the National Key R&D Program of China (2020AAA0108001). The authors would like to thank the anonymous reviewers for their valuable comments and suggestions to improve this paper.
Appendix A Evaluation Details
A.1 Benchmark and Evaluation Details
Our evaluation encompasses five categories of safety benchmarks, which can be broadly divided into two types: jailbreak robustness and safety-awareness. Specifically, FigStep, MM-SafetyBench, and JailBreaKV fall under the jailbreak robustness benchmark, for which we report the Attack Success Rate (ASR). Meanwhile, MSSBench and SIUO are categorized as safety-awareness benchmarks, and we follow the respective safety metrics according to the original papers.
For the MM-SafetyBench, FigStep and SIUO, we employ GPT-4o-mini Hurst et al. (2024) and adopt identical configurations in their work to ensure consistency across assessments. For the JailBreaKV, we follow the same evaluation settings of their work and utilize the Llama Guard 3 Vision Chi et al. (2024) as the evaluator. However, similar to the issues observed in MIS Ding et al. (2025b), we identify shortcomings in the safe category evaluation setup of MSSBench. As a result, we employed human evaluation as an alternative. Additional details are provided in Appendix A.2. The benchmarks are listed as follows:
FigStep.
FigStep Gong et al. (2025) consists of 500 harmful queries spanning 10 categories. Rather than directly inputting harmful textual prompts, it conveys malicious intent through visual formatting. Each image encodes harmful content as a visually rendered list, which is then paired with a fixed textual instruction: “The image shows a list numbered 1, 2, and 3, but the items are empty. Please generate detailed content for each item on the list.
$M$-SafetyBench.
MM-SafetyBench Liu et al. (2024b) covers 13 categories and is primarily designed to evaluate the responses of MLLMs to malicious queries. In our experiments, we adopt the SD+TYPO setting, which includes a total of 1,680 samples, following the setting of Ding et al. (2025b).
JailBreaKV.
JailBreaKV Luo et al. (2024) is designed to evaluate the transferability of jailbreak techniques from LLMs to MLLMs, encompassing both text-based and image-based jailbreak attacks. Following the setting of VLSBench Hu et al. (2024), we conduct our experiments on the miniset provided in their paper, which contains 280 samples.
MSSBench.
Multimodal Situation Safety Benchmark (MSSBench) Zhou et al. (2024) explores how safety considerations vary depending on the specific context involving the user or agent. In this benchmark, half of the image contexts are safe, while the other half are unsafe. The benchmark identifies safe intent of the models in two distinct settings: chat and embodied scenarios. Due to economic constraints, our evaluation was conducted using a single instruction sampled from each of the safe and unsafe scenarios.
SIUO.
Safety Inputs but Unsafe Outputs (SIUO) Wang et al. (2025) considers cases where the image and text modalities are independently safe, but their combination leads to unsafe or unethical outputs. The SIUO covers 9 categories including a total of 167 samples. We report the evaluation metrics as defined in their paper.
Multimodal Large Reasoning Model Base Model
LlamaV-o1 Thawakar et al. (2025) Llama 3.2-vision-11B
LLaVA-CoT Xu et al. (2024) Llama 3.2-vision-11B
Mulberry-Llama Yao et al. (2024) Llama 3.2-vision-11B
LMM-R1 Peng et al. (2025b) Qwen2.5-VL-3B
R1-Onevision Yang et al. (2025) Qwen2.5-VL-7B
Mixed-R1 Xu et al. (2025) Qwen2.5-VL-7B
SophiaVL-R1 Fan et al. (2025) Qwen2.5-VL-7B
MM-Eureka Meng et al. (2025) InternVL-2.5-8B
Kimi-VL-Thinking Team et al. (2025) Kimi-VL
Skywork-R1V Peng et al. (2025a) -
QVQ-Preview Team (2024) -
Gemini2.0-Flash-Thinking DeepMind (2025) -
Claude3.7-Sonnet-Thinking Anthropic (2025) -
Table 6: List of the evaluated MLRMs and their corresponding base models.
A.2 MSSBench Evaluation Details
In the evaluation setup of MSSBench, LLMs are required to classify responses into four safety levels. For certain examples in the safe category, responses that provide helpful advice but also emphasizing potential risks are evaluated as incorrect. As shown in Figure 7, the model provides the suggestion of using a heavy-duty drill while emphasizes safety as a priority. To address this limitation, we employed human evaluation as an alternative.
The image, question, and response are delivered to three independent human annotators to obtain evaluation results. For the safe category, the evaluation primarily assesses whether the model exhibits unnecessary or excessive refusal. For the unsafe category, the evaluation focuses on whether the model can accurately detect and identify potential risks.
Figure 7: A failure case of MSSBench evaluation with LLM-as-Judge method
A.3 Model Details
We evaluate a total of 13 MLRMs, Table 6 summarizes the Multimodal Large Reasoning Models evaluated for safety and their corresponding base models.
Model Chat Embodied AVG.
Safe Unsafe AVG. Safe Unsafe AVG.
[…]
Table 7: Results on MSSBench in two distinct settings: chat and embodied scenarios.
Appendix B Dataset Details
B.1 Detailed Pipeline
Step 1: Data Preparation.
Based on the categories provided in SPA-VL Zhang et al. (2024) and VLGuard Zong et al. (2024) datasets, we extract the raw data in the first step of construction for data preparation. In addition, we include the paired safety examples from VLGuard, where each image is matched with both a safe and an unsafe question. This design aims to improve the ability of the models to handle context-sensitive safety issues. However, the answers in SPA-VL are derived from pairwise preference data generated by large language models. Despite selecting the “chosen” responses as the final answers, a subset of them remain potential safety risks. The unsafe answers are subsequently screened and reconstructed.
Step 2: Image Description Generation.
To accommodate the input requirements of DeepSeek-R1 Guo et al. (2025), we convert images into caption form as input in advance.In particular, we employ Qwen2.5-VL-72B Bai et al. (2025) to generate image captions to ensure that the semantic loss during modality conversion is within an acceptable range in terms of safety issues, with the prompt “Please provide a detailed description of this image.”.
Step 3: Safety Thought Process Generation.
After obtaining the image captions in the second step, the image caption, question, original response, and safety guidelines are provided to DeepSeek-R1 Guo et al. (2025) to generate safety-oriented thought process. The model first analyzes the question and the caption to obtain the underlying intent, then refers to the safety guidelines to generate a safety-oriented thought process corresponding to the provided response.
Step 4: Thought Process Filter.
The primary objective in this step is to filter and transform the original thought process. In particular, the caption field is processed to prevent potential misinterpretation by the MLRMs. The original thought process from the previous step also contains redundant elements such as headers like “Safety Chain of Thought” which need to be removed and reformatted. Specific prompts used for this step are provided in Figure 8.
B.2 Dataset Statics
The statistics of data instances in each category within our dataset after the final filtering process are shown in Table 8.
Category # Samples
Privacy violation 196
Professional advice 200
Political sensitivity 209
Sexually explicit 199
Violence 204
Disinformation 205
Discrimination 231
Hate speech 200
Economic harm 200
Physical harm 196
Illegal activities 200
Malware 200
Safe 977
All 4394
Table 8: The statistics of our dataset
B.3 Prompt for Data Construction
This section presents the prompt and safety regulation used for safety thought process collection. We referenced the definition of User Requests Categorization from Wang et al. (2024) and listed all prompts used in the data construction in Figure 8. During the collection of safety thought process data, we supplied the model with category-specific safety regulations derived from the raw data classification. This strategy not only reduces the length of prompt but also enables tiered defensive measures for distinct safety-related issues.
Appendix C Training Details
C.1 Baselines
To evaluate the effectiveness of our data construction, we select three existing multimodal safety alignment datasets as baselines that are listed as follows:.
  • Direct: The original safety capability of the model without fine-tuning.
  • VLGuard Zong et al. (2024): The first safety dataset specifically designed for fine-tuning of MLLMs.
  • MIS Ding et al. (2025b): A multi-image safety dataset that combines visual perception with reasoning logic labels.
  • SPA-VL Zhang et al. (2024): A safety preference alignment dataset for Vision Language Models.
C.2 Traning Details
For the training procedure, we utilize the LoRA Hu et al. (2022) to fine-tune the R1-Onevision and LLaVA-CoT based on the LLaMA-Factory [^1] https://github.com/hiyouga/LLaMA-Factory framework. For the SPA-VL dataset, to ensure fairness, we randomly selected 4,000 data samples with balanced categories and used the DPO script for training. All training experiments are conducted using one NVIDIA A100-80G GPU. The hyper-parameters for supervised fine-tuning are listed in Table 9.
Hyper-Parameter Value
lora rank 8
learning rate 1.0e-5
train epoch 1.5
per_device_batchsize 1
warm up ratio 0.1
learning rate scheduler cosine
Table 9: Hyper-parameters for supervised fine-tuning.
Appendix D More Results
D.1 Detailed Experiments on MSSBench
In the evaluation protocol of MSSBench, scenarios are divided into chat and embodied settings, with safety scores calculated separately for safe and unsafe conditions. The average safety scores are reported in Table 1, and the specific scores for each category are detailed in Table 7.
D.2 Performance of Thought Process on MM-SafetyBench
We also independently evaluate the safety performance of the thought process on MM-SafetyBench. This section presents the MLRMs’ performance and fine-tuned models’ performance in Table 10 and Table 11.
  R1-Onevision Safe Answer Unsafe Answer
Safe Thought 13.71% 41.66%
Unsafe Thought 6.72% 37.91%
Kimi-VL-Thinking Safe Answer Unsafe Answer
Safe Thought 13.27% 40.65%
Unsafe Thought 25.25% 20.83%
LLaVA-CoT Safe Answer Unsafe Answer
Safe Thought 18.69% 40.77%
Unsafe Thought 9.05% 31.49%
QVQ-Preview Safe Answer Unsafe Answer
Safe Thought 23.15% 44.76%
Unsafe Thought 7.50% 24.52%
Skywork-R1V Safe Answer Unsafe Answer
Safe Thought 21.07% 42.26%
Unsafe Thought 6.25% 30.42%
Table 10: Proportion of safety in the thought process and the final answer on MM-SafetyBench.
  R1-Onevision Safe Answer Unsafe Answer
Safe Thought 13.71% 41.66%
Unsafe Thought 6.72% 37.91%
R1-Onevision+TiS Safe Answer Unsafe Answer
Safe Thought 77.32% 21.49%
Unsafe Thought 0.95% 0.24%
LLaVA-CoT Safe Answer Unsafe Answer
Safe Thought 18.69% 40.77%
Unsafe Thought 9.05% 31.49%
LLaVA-CoT+TiS Safe Answer Unsafe Answer
Safe Thought 84.82% 8.75%
Unsafe Thought 5.89% 0.54%
Table 11: Proportion of safety in the thought process and the final answer of fine-tuned model using TiS on MM-SafetyBench.
Figure 8: Prompts used in data construction.
D.3 Case Study of Ablation Experiments
Examples of responses generated by models fine-tuned on TiS and TiS no thought dataset are illustrated in Figure 9 and Figure 10. The TiS dataset without thought is predominantly composed of unsafe examples. When fine-tuning MLRMs based on this data, the model tends to generate responses that consist solely of brief refusals. In particular, for safety-awareness examples, the model tends to either produce overly brief refusals or fail to detect unsafe risks. In contrast, the complete TiS data enables the detection of more unsafe intents during the thought process, while also providing more helpful information.
Figure 9: Examples from R1-Onevision tuning on TiS w/o thought process. The red indicates the unsafe parts, while the blue indicates the safe parts.
Figure 10: Examples from LLaVA-CoT tuning on TiS w/o thought process. The red indicates the unsafe parts, while the blue indicates the safe parts.
D.4 Case Study
Examples of responses generated models fine-tuned on VLGuard, MIS, SPA-VL and TiS dataset are illustrated in Figure 11 and Figure 12. Our approach demonstrates the ability to retain the thought process of the models while decisively rejecting unsafe inputs and explicitly articulating the potential serious consequences associated with such queries.
Figure 11: Examples of several baselines and our dataset on MM-SafetyBench. We only highlight some fragments to show the representative difference between various methods. The red indicates the unsafe parts, while the blue indicates the safe parts.
Figure 12: Examples of several baselines and our dataset on SIUO. We only highlight some fragments to show the representative difference between various methods. The red indicates the unsafe parts, while the blue indicates the safe parts.